The point (-5+bi) has been plotted on the complex plane and r=9. If 0°≤θ≤180° find the value of b. Leave your answer as a radical.

Thank you all for your time and assistance!
-ClanMacniall

Answer by stanbon(75887)

The point (-5+bi) has been plotted on the complex plane and r=9. If 0°≤θ≤180° find the value of b. Leave your answer as a radical.
-----------
With that range for theta, (-5+bi) must be in the 2nd Quadrant
and b must be positive.
----
x = -5 and r = 9
---
Solve for "y"
9^2 = [5^2+y^2]
81 = 25 + y^2
y^2 = 56
y = sqrt(56)
----
b = y = sqrt(56)
===================
Cheers,
Stan H.